# An Empirical Evaluation of Pre-trained Large Language Models for Repairing Declarative Formal Specifications

### v1

@@ 5. Experimental Results
5.1. Results for RQ1: Effectiveness
@@
Table 5. Comparing against state-of-the-art Alloy repair techniques (ARepair (Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), ATR (Zheng et al., 2022), and Hasan et al. (Hasan et al., 2023)) on the ARepair and Alloy4Fun benchmarks, using GPT-3.5-Turbo as our LLM baseline.
Total Arepair ICEBAR BeAFix ATR Hasan et al. No-Feedback Generic-Feedback Auto-Feedback
#specs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs
Alloy4Fun
1936 185 1051 981 1286 147 156 231 278
ARepair
38 9 21 24 22 4 4 6 18
Table 6. Comparing with state-of-the-art Alloy repair techniques (ARepair (Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), ATR (Zheng et al., 2022), and Hasan et al. (Hasan et al., 2023)) on the ARepair benchmarks, using various combinations of LLMs and report levels as outlined in Table 4.
@@
Model Total Arepair ICEBAR BeAFix ATR Hasan et al. Setting-1 Setting-2 Setting-3 Setting-4 Setting-5 Setting-6
#specs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs
ARepair addr 1 1 1 1 1 0 0 0 1 0 1 1
arr 2 2 2 2 1 0 0 0 0 1 0 1
balancedBST 3 1 2 1 1 1 1 1 2 1 1 1
bempl 1 0 1 0 1 0 0 0 0 1 1 1
cd 2 0 2 2 2 1 1 2 2 2 1 2
ctree 1 1 0 0 0 1 1 1 1 1 1 1
dll 4 0 3 3 2 4 4 4 4 2 2 4
farmer 1 0 0 0 0 1 1 1 1 1 1 1
fsm 2 2 2 1 2 1 1 1 2 0 0 1
grade 1 0 1 0 1 0 0 0 0 0 0 1
other 1 0 0 1 1 0 0 0 0 0 0 1
student 19 2 7 13 10 6 6 6 9 8 11 13
Summary 38 9 21 24 22 15 15 16 22 17 19 28
Figure 2. Venn diagrams illustrating the exclusive and overlapping cases of successful repairs by different repair methods, highlighting the complementary and unique capabilities of the APR pipeline compared to SoTA repair tools.
Table 5 provides a comparison using GPT-3.5-Turbo as the baseline. The column ”Total #specs” indicates the total number of defective specifications in each benchmark, while the subsequent columns show the number of correctly repaired specifications by each state-of-the-art Alloy repair tool. The last three columns display the repair outcomes based on different report levels employed in the pipeline. The performance of the baseline LLM (GPT-3.5-Turbo) is notably inferior compared to traditional state-of-the-art tools, consistent with prior findings by Hasan et al. (Hasan et al., 2023). However, the two-agent (auto-feedback) configuration demonstrated the most promising performance among the feedback levels across both benchmarks.
Table 6 presents comparison results for the ARepair benchmark using GPT-4-32k and GPT-4-Turbo. The ”Models” column categorizes model groups in the ARepair benchmark, while the last six columns illustrate the total count of accurately fixed specifications under different pipeline settings (summarized in Table 4). Setting-6 notably outperforms all state-of-the-art tools, while Setting-3 performs similarly to BeAFix and surpasses other traditional tools. It’s noteworthy that the specification farmer remained unresolved by all SoTA repair tools, but was successfully repaired in the APR pipeline across all settings. This suggests that GPT-4 family models can effectively address complex specifications, particularly those with multi-line bugs. Further, the results observed in Setting-3 and Setting-6 surpass those of other settings (in Table 4) and align with the ”Auto-Feedback” column in Table 5. This alignment underscores the efficacy of prompt construction using LLMs over generic prompts developed by humans, advocating for the use of multi-agent LLM setups.
To gain deeper insights into the complementarity of these methods, Venn diagrams in Figure 2 illustrate cases exclusively fixed by one tool but not by others, as well as the overlap between techniques (i.e., cases correctly repaired by multiple tools). Notably, the repair pipeline successfully addresses faulty models that other SoTA repair tools fail to resolve.
The metric Correct@6 is utilized to compare the performance of all LLMs across various feedback levels. Table 7 showcases the comparison results using the ARepair benchmark. While GPT-3.5-Turbo exhibited less effective repair capabilities, GPT-4-Turbo demonstrated superior performance across different feedback levels, attributable to its cut-off and context window parameters relative to its counterparts (see Table 3), alongside enhanced features integrated into this model (detailed in Section 4.4).
The results for RQ1 demonstrate that the APR pipeline, particularly with GPT-4-Turbo and multi-agent setups, outperforms traditional state-of-the-art Alloy repair tools, effectively addressing complex multi-line bugs and showcasing superior performance across various feedback levels.
@@
Table 7. Comparison results presenting the repair performance of different LLMs across various feedback levels, i.e., No-Feedback (NF), Generic-Feedback (GF), Auto-Feedback (AF), measured by the Correct@6 metric.
GPT-3.5-Turbo GPT-4-32k GPT-4-Turbo
NF GF AF NF GF AF NF GF AF
10.5 15.8 47.4 39.5 42.1 57.9 44.7 50.0 57.9
5.2. Results for RQ2: Performance
This research question delves into the repair performance of various pre-trained Language Model Models (LLMs) concerning factors such as repair iteration budget, bug type, localization capabilities, and adherence to instructions. Our analysis focuses on two LLM variants: GPT-4-32k and GPT-4-Turbo, utilizing the results provided in Table 6 and the settings outlined in Table 4.
@@ 5.2. Results for RQ2: Performance
Repair Iteration Budget: The box-whisker plot in Figure 3(a) illustrates the iterations needed by different settings to rectify faulty specifications. Across all settings, the median number of iterations predominantly centers around 1.0, except for Setting 3 & 6, where Setting-3 increases to 2.5, indicating a demand for additional iterations. However, median values for Settings 1, 2, 4, and 5 remain at 1.0, suggesting that at least half of the issues are resolved within the first iteration.
@@
Figure 3. (a) Iteration count distribution for repairing specifications across various settings. (b) Repair distribution in ARepair benchmark categorized by bug type: single-line or multi-line (higher values preferred). (c) The incident count for initial repair attempts mirroring the buggy specification (lower values preferred), derived from the ARepair benchmark.
Repairing Based on Bug Type: Figure 3(b) depicts the repair capabilities of GPT-4-32k and GPT-4-Turbo concerning bug types, namely single-line and multi-line bugs. The best performance for GPT-4-Turbo (auto-feedback, Setting-6) achieved the highest repair rate, fixing 16 single-line bugs and 12 multi-line bugs. In contrast, the best result for GPT-4-32k (auto-feedback, Setting-3) demonstrated superior performance in correcting single-line bugs (12 models) compared to multi-line bugs (10 models). Additionally, we examine the repair outcomes achieved by GPT-3.5-Turbo, showing repair capabilities for both single-line and multi-line bugs.
@@
Following Instructions: Adherence to instructions is crucial in tasks like APR to ensure repair quality and control costs. However, Figure 3(c) reveals that both GPT-4 models sometimes repeat buggy models in their initial repair iterations, despite system instructions to avoid this. GPT-4-Turbo generally shows better compliance, particularly in Setting-6, where it exhibits the lowest number of repeated buggy specifications.
The comparison between the APR pipeline utilizing different feedback levels and state-of-the-art Alloy repair techniques, along with the performance of various LLMs, revealed that the auto-feedback approach exhibited the best performance across both benchmarks, surpassing traditional SoTA repair tools.
5.3. Results for RQ3: Adaptive Prompting
The APR pipeline facilitates three feedback levels—No-feedback, Generic-Feedback, and Auto-Feedback—as outlined in Section 2. This research question delves into the implications of these feedback levels on LLM performance, drawing insights from the following analysis:
Overall Repair Performance: Examination of Tables 5 and 6 reveals a notable correlation between the use of ”auto-feedback” and superior repair efficacy across all LLMs, including GPT-3.5-Turbo. Conversely, the absence of feedback, characterized by ”No-feedback,” resulted in the poorest repair performance, indicating that prompt updates driven by information from the Alloy Analyzer report significantly contribute to repair success.
Bug-type Repair: Across all LLMs, Settings 3 and 6 consistently demonstrate a higher ability to repair complex specifications, as illustrated in Figure 3(b). This suggests that adaptive prompting, particularly in Settings 3 and 6, enhances the pipeline’s capacity to address intricate bugs effectively.
Iteration Budget Analysis: Figure 3(a) showcases a slight increase in the requirement for additional iterations in Settings 3 and 6. While these settings exhibit superior repair capabilities, they may necessitate a marginally higher number of iterations to achieve optimal outcomes.
These observations underscore the advantage of leveraging LLMs to automatically generate prompts based on Alloy Analyzer reports, leading to enhanced repair performance compared to Generic-Feedback. However, this advantage may be accompanied by the need for additional iterations.
5.4. Results for RQ4: Characteristics of Failed Repairs
This research inquiry delves into the root causes of failures encountered during the iterative repair process across various settings. The following elucidates the key findings derived from this investigation.
Figure 4. Error types observed in failed repair iterations across all settings

### v2

@@ 5 Experimental Results
5.1 Results for RQ1: Effectiveness
This research question investigates the effectiveness of LLMs in repairing Alloy specifications under various settings within the repair pipeline. To address this, we (1) compare the repair pipeline with state-of-the-art Alloy repair tools and (2) assess the impact of different feedback levels, LLMs, and agent paradigms on repair performance, and study the influence of LLMs on the repaired models.
Comparing with SoTA. Tables 5 and 6 present a comprehensive comparison of our approach against state-of-the-art Alloy repair techniques, including ARepair Wang et al. (2019), ICEBAR Gutiérrez Brida et al. (2023), BeAFix Brida et al. (2021), ATR Zheng et al. (2022), and Hasan et al. Hasan et al. (2023), on the ARepair and Alloy4Fun benchmarks, respectively. Each table reports the number of defects addressed and repairs achieved by each tool, as well as by various LLM settings described in Table 4. For the ARepair benchmark (Table 5), results are shown across all settings, enabling a detailed evaluation of repair effectiveness. For the Alloy4Fun benchmark (Table 6), GPT-4o is applied to a representative subset of models to balance performance and computational cost, while GPT-3.5-Turbo is evaluated on the full dataset.
The “Defects Count” column reports the total number of bugs identified within each model category, providing insight into the distribution of defects across specifications. When the value in “Defects Count” is equal to “Total #specs,” it indicates that each specification contains exactly one defect, suggesting the presence of only single-line bugs. In contrast, if “Defects Count” exceeds “Total #specs,” this reveals that some specifications contain multiple defects, indicative of multi-line bugs. For example, in Table 5, the balancedBST model exhibits 8 defects across 3 specifications, clearly demonstrating the occurrence of multi-line bugs.
The subsequent columns present the number of correctly repaired specifications for each state-of-the-art Alloy repair tool, with the remaining columns reporting the repair outcomes under the various settings described in Table 4.
The results highlight the superior performance of the GPT-4 family—particularly GPT-4o and GPT-4-Turbo—over both GPT-4-32k and traditional state-of-the-art Alloy repair tools. In contrast, GPT-3.5-Turbo (Settings 10–12) demonstrates significantly lower repair effectiveness than traditional tools, which aligns with previous findings by Hasan et al. Hasan et al. (2023).
Furthermore, across all evaluated LLMs and benchmarks, the Auto-feedback configuration (Settings 3, 6, 9, and 12 in Tables 6 and 5) consistently delivers the strongest repair performance. Notably, the Generic-feedback configuration (Setting 8) achieves results comparable to Setting 9 on the Alloy4Fun sampled benchmark with GPT-4o, a phenomenon attributed to the inherent stochasticity of LLMs. To further investigate, we re-executed repair experiments on five models that were initially only repaired under Generic-feedback; two of these were successfully repaired upon rerunning with Auto-feedback. This outcome underscores the effectiveness of LLM-driven prompt construction over generic, human-crafted prompts and supports the adoption of multi-agent LLM configurations for automated specification repair.
@@ 5.1 Results for RQ1: Effectiveness
Table 5: Comparison of state-of-the-art Alloy repair techniques—ARepair Wang et al. (2019), ICEBAR Gutiérrez Brida et al. (2023), BeAFix Brida et al. (2021), ATR Zheng et al. (2022), and Hasan et al. Hasan et al. (2023)—on the ARepair benchmark, using the settings detailed in Table 4.
@@
Model Total Defects ARepair ICEBAR BeAFix ATR Hasan et al. #repairs Under Setting
#specs Count #repairs #repairs #repairs #repairs #repairs GPT-4-32k GPT-4-Turbo GPT-4o GPT-3.5-Turbo
1 2 3 4 5 6 7 8 9 10 11 12
addr 1 1 1 1 1 1 0 0 0 1 0 1 1 0 0 0 0 0 0
arr 2 2 2 2 1 0 0 0 0 0 1 0 1 1 2 1 0 1 1
balancedBST 3 8 1 2 1 1 0 1 1 2 1 1 1 1 1 1 0 0 1
bempl 1 1 0 1 0 1 1 0 0 0 1 1 1 0 0 0 0 0 0
cd 2 3 0 2 2 2 0 1 2 2 2 1 2 2 2 2 1 0 0
ctree 1 1 1 0 0 0 0 1 1 1 1 1 1 1 1 1 0 0 0
dll 4 8 0 3 3 2 0 4 4 4 2 2 4 4 4 4 1 3 3
farmer 1 2 0 0 0 0 0 1 1 1 1 1 1 1 1 1 1 1 1
fsm 2 2 2 2 1 2 0 1 1 2 0 0 1 1 1 2 0 0 0
grade 1 1 0 1 0 1 0 0 0 0 0 0 1 0 0 1 0 0 0
other 1 1 0 0 1 1 1 0 0 0 0 0 1 1 1 1 1 0 1
student 19 34 2 7 13 10 2 6 6 9 8 11 13 12 14 14 0 1 11
Summary 38 64 9 21 24 22 4 15 16 22 17 19 28 24 27 28 4 6 18
%repair 23.7 55.3 63.2 57.8 10.5 39.5 42.1 57.9 44.7 50 73.4 63.2 71.1 73.4 10.5 15.8 47.4
Table 6: Comparison of state-of-the-art Alloy repair techniques—ARepair Wang et al. (2019), ICEBAR Gutiérrez Brida et al. (2023), BeAFix Brida et al. (2021), ATR Zheng et al. (2022), and Hasan et al. Hasan et al. (2023)—on the Alloy4Fun benchmark, using the settings described in Table 4. Bracketed numbers indicate the models selected for the GPT-4o repair experiment to manage computational costs.
Model Total Defects ARepair ICEBAR BeAFix ATR Hasan et al. #repairs Under Setting
#specs Count #specs #repairs #repairs #repairs #repairs GPT-4o GPT-3.5-Turbo
7 8 9 10 11 12
classroom 999(60) 999(60) 88(3) 424(18) 387(16) 688(29) 88(4) (0) (0) (6) 0 1 21
cv 138(60) 138(60) 2(0) 86(26) 82(20) 38(0) 4(3) (56) (58) (59) 38 32 70
graphs 283(60) 283(60) 19(8) 237(39) 232(38) 260(44) 20(39) (22) (40) (22) 23 51 35
lts 249(60) 249(60) 1(0) 73(20) 41(9) 70(15) 21(32) (38) (48) (46) 68 115 106
production 61(60) 61(60) 27(23) 36(33) 56(53) 43(46) 12(12) (59) (60) (58) 25 32 34
trash 206(57) 206(57) 48(9) 195(46) 183(49) 187(34) 2(2) (1) (0) (10) 0 0 12
Summary 1936(357) 1936(357) 185(43) 1051(182) 981(185) 1286(168) 147(92) (176) (206) (201) 154 231 278
%repair 9.5(12) 54.2(50.9) 50.6(51.8) 66.4(47) 7.6(25.7) (49.2) (55.7) (56.3) 7.6 11.9 14.3
The repair results for the ARepair benchmark, as presented in Table 5, show that Settings 6 and 9 deliver substantially superior performance compared to all SoTA tools, including traditional program repair approaches. Notably, the APR pipeline is able to repair specifications that remain unresolved by most, or even all, existing SoTA techniques. For instance, the grade specification (see Listing 1), which could not be repaired by the majority of SoTA tools, was successfully addressed under both Settings 6 and 9. Similarly, the farmer specification, which contains a multi-line bug and was not repaired by any SoTA tool, was successfully fixed by all APR configurations. The ctree specification, which was addressed by only a single SoTA tool, was also repaired by all APR configurations leveraging GPT-4 models.
These results demonstrate that models from the GPT-4 family are particularly effective at handling complex specifications, especially those involving multi-line bugs. In particular, the findings underscore the promising potential of agentic LLMs enhanced with feedback mechanisms. As shown in Table 7, which provides a comparative analysis of repair performance across all evaluated LLMs using the Correct@6 metric, the dual-agent paradigm (Auto-Feedback) consistently outperforms the single-agent approach across all models. Within this paradigm, GPT-4o achieved the highest repair performance.
@@
Table 7: Comparison of repair performance for different LLMs across various feedback levels—No-Feedback (NF), Generic-Feedback (GF), and Auto-Feedback (AF)—as measured by the Correct@6 metric.
GPT-3.5-Turbo GPT-4-32k GPT-4-Turbo GPT-4o
Single-agent Dual-agent Single-agent Dual-agent Single-agent Dual-agent Single-agent Dual-agent
NF GF AF NF GF AF NF GF AF NF GF AF
10.5 15.8 47.4 39.5 42.1 57.9 44.7 50.0 73.4 63.2 71.1 73.4
Table 8: Average number of variables and clauses in propositional formulas generated by the Alloy Analyzer from LLM-produced repairs on the ARepair benchmark. Results are reported as overall averages across all evaluated settings for each LLM model.
Model Total # spec LLM Models
GPT-4-32k GPT-4-Turbo GPT-4o GPT-3.5-Turbo
addr 1 variables: 333 variables: 567 variables: 66 variables: 187
clauses: 368 clauses: 627 clauses: 74 clauses: 215
arr 2 variable: 1,973 variable: 7,578 variables: 6,364 variables: 1,482
clauses: 4,969 clauses: 16,284 clauses: 6,096 clauses: 3,314
balancedBSt 3 variables: 2,001 variables: 15,312 variables: 4,155 variables: 1,797
clauses: 3,720 clauses: 52,602 clauses: 8,616 clauses: 3,306
bempl 1 variables: 2,249 variables: 700 variables: 744 variables: 694
clauses: 3,489 clauses: 1,012 clauses: 1,071 clauses: 1,003
cd 2 variables: 294 variables: 280 variables: 288 variables: 576
clauses: 196 clauses: 164 clauses: 326 clauses: 1,188
ctree 1 variables: 2,976 variables: 1,130 variables: 1,721 variables: 1,679
clauses: 2,857 clauses: 1,035 clauses: 1,591 clauses: 1,524
dll 4 variables: 21,572 variables: 20,012 variables: 21,524 variables: 6,324
clauses: 66,088 clauses: 65,248 clauses: 50,128 clauses: 4,108
farmer 1 variables: 1,468 variables: 1,213 variables: 2,252 variables: 1,199
clauses: 2,808 clauses: 2,177 clauses: 4,096 clauses: 2,429
fsm 2 variables: 894 variables: 2,224 variables: 540 variables: 1,080
clauses: 1,816 clauses: 4,556 clauses: 1,108 clauses: 2,278
grade 1 variables: 4,763 variables: 2,830 variables: 1,024 variables: 491
clauses: 7,930 clauses: 4,578 clauses: 1,585 clauses: 604
other 1 variables: 477 variables: 264 variables: 113 variables: 268
clauses: 78 clauses: 352 clauses: 128 clauses: 358
Student 19 variables: 100,548 variables: 63,992 variables: 31,236 variables: 63,878
clauses: 321,727 clauses: 208,696 clauses: 20,520 clauses: 208,354
Total 38 variables: 139,548 variables: 116,102 variables: 70,027 variables: 79,655
clauses: 416,046 clauses: 357,331 clauses: 95,339 clauses: 228,681
Table 9: Average number of variables and clauses in propositional formulas generated by the Alloy Analyzer from LLM-produced repairs on the A4F benchmark. Results are reported as overall averages across all evaluated settings for each LLM model. For a fair comparison, results are based on the same set of 357 specifications for both GPT-3.5-Turbo and GPT-4o.
GPT-4o GPT-3.5-Turbo
Model # specs used Variables # specs used Variables
classroom 60 variables: 35,760 60 variables: 3,580
clauses: 47,340 clauses: 4,691
cv 60 variables: 105,660 60 variables: 31,320
clauses: 8,640 clauses: 34,620
graphs 60 variables: 17,040 60 variables: 7,902
clauses: 37,200 clauses: 6,180
lts 60 variables: 26,100 60 variables: 88,560
clauses: 26,820 clauses: 199,260
production 60 variables: 16,560 60 variables: 61,380
clauses: 19,560 clauses: 124,560
trash 57 variables: 7,068 57 variables: 23,598
clauses: 8,892 clauses: 43,605
Total 357 variables: 208,188 357 variables: 216,340
clauses: 148,452 clauses: 412,916
To gain deeper insights into the complementarity of these methods, the Venn diagrams in Figure 2 illustrate cases fixed exclusively by one tool but not by others, as well as the overlap between techniques (i.e., cases correctly repaired by multiple tools). In particular, the repair pipeline successfully addresses faulty models that other SoTA repair tools cannot resolve.
Figure 2: Venn diagrams showing the exclusive and overlapping successful repairs for the ARepair benchmark achieved by different repair methods. These illustrate the complementary and unique capabilities of the APR pipeline compared to state-of-the-art (SoTA) repair tools.
Influence of Defective Model Complexity/Size on Repair Performance. To investigate the relationship between model complexity and repair performance, we assessed complexity based on the number of variables clauses in the propositional formulas generated by the Alloy Analyzer from LLM-produced repairs (see Tables 8 and 9) and conducted a detailed analysis thereof.
GPT-4o consistently outperforms GPT-3.5-Turbo in repairing faulty specifications, despite the latter generating more complex formulas with significantly more clauses. The clause-to-variable ratio further highlights this difference (2.87 for GPT-3.5-Turbo vs. 1.36 for GPT-4o), indicating that GPT-4o constructs more concise and efficient logical expressions. This suggests that GPT-4o achieves more efficient and accurate repairs by leveraging simpler logical structures while maintaining correctness.
The tendency of GPT-3.5-Turbo to generate more verbose and constraint-heavy formulas, even with comparable variable counts, underscores the limitations imposed by its smaller context window and less sophisticated reasoning capabilities. The consistently higher clause-to-variable ratios across benchmarks for GPT-3.5-Turbo further support this conclusion. These findings confirm that model capability—especially in context comprehension and logical reasoning—is a critical factor for successful Alloy repair. GPT-4o’s superior performance, as highlighted in the %repair row of Table 6, demonstrates how advances in LLM architecture directly enhance formal reasoning and model repair effectiveness.
Discussing LLMs repair performance. The evaluation reveals a clear performance gap between GPT-3.5-Turbo and the GPT-4 family across all experimental configurations. GPT-3.5-Turbo consistently exhibits the lowest repair rates, largely due to its architectural limitations such as a smaller context window and less advanced reasoning capabilities. These constraints hinder its ability to effectively incorporate iterative feedback and maintain coherence over successive repair attempts, often causing context overflows and semantic inconsistencies. Additionally, GPT-3.5-Turbo frequently generates Alloy-specific errors, including incorrect operator usage, type mismatches, and hallucinated constructs, that are less common in specialized repair tools designed with native Alloy semantics. These shortcomings persist even in simpler Alloy4Fun models, limiting its practical utility in automated specification repair. In contrast, GPT-4 models, including GPT-4-32k, GPT-4-Turbo, and GPT-4o, demonstrate significantly improved repair performance across all settings and benchmarks. Benefiting from larger context windows and enhanced reasoning abilities, these models handle complex specifications and multi-line bugs more effectively. Among them, GPT-4o and GPT-4-Turbo achieve comparable results, with GPT-4o often leading in agentic (Auto-Feedback) configurations that leverage multi-agent prompt construction and iterative feedback. This dual-agent paradigm consistently boosts repair success, highlighting the advantages of advanced LLM architectures combined with adaptive prompting strategies.
The results for RQ1 indicate that an agentic LLM incorporating an iterative feedback mechanism achieves improved repair performance compared to both traditional state-of-the-art Alloy repair tools and non-agentic LLM configurations. Notably, GPT-4o consistently outperformed other LLMs, yielding the most effective repair outcomes across both agentic and non-agentic settings.
5.2 Results for RQ2: Performance
This research question delves into the repair performance of various pre-trained LLMs concerning factors such as repair iteration budget, bug type, localization capabilities, and adherence to instructions. This analysis leverages a comprehensive benchmark that includes all evaluated LLMs and a diverse range of defect types, encompassing both single-line and multi-line bugs. As shown in Table 5, each defective model was tested across up to 12 distinct settings and up to 6 repair iterations, resulting in a total of 2,736 repair trials—effectively corresponding to the repair of 2,736 defective models. This extensive experimental setup provides a robust and representative evaluation of the repair capabilities of the LLMs across a wide variety of scenarios.
@@ 5.2 Results for RQ2: Performance
Repair Iteration Budget. The box-whisker plot in Figure 3(a) illustrates the iterations needed by different settings to rectify faulty specifications. Across all settings, the median number of iterations predominantly centers around 1.0, except for Settings 3, 6, and 10-12. Although Settings 3 and 6 exhibit superior repair capabilities, they may require a marginally higher number of iterations to achieve optimal results. However, median values for Settings 1, 2, 4, and 5 remain at 1.0, suggesting that at least half of the issues are resolved within the first iteration. In contrast, the median values for Settings 7–9 are consistently 1.0, indicating that such settings require the fewest repair iterations, independent of the employed feedback mechanism.
@@
Figure 3: (a) Iteration count distribution for repairing specifications across various settings; (b) Repair distribution categorized by bug type: single-line or multi-line (higher values preferred); (c) The incident count for initial repair attempts mirroring the buggy specification (lower values preferred).
Repairing Based on Bug Type. Figure 3(b) illustrates the repair effectiveness of various configurations in the repair pipeline with respect to bug types—specifically, single-line and multi-line defects. The dual-agent configurations in Setting-6 (auto-feedback, GPT-4-Turbo) and Setting-9 (auto-feedback, GPT-4o) demonstrated the highest overall repair rates, successfully fixing 16 single-line and 12 multi-line bugs, and 15 single-line and 13 multi-line bugs, respectively. In comparison, the best-performing configuration for GPT-4-32k (Setting-3, auto-feedback) showed relatively stronger performance on single-line bugs, repairing 12 Alloy models, versus 10 for multi-line bugs.
@@
Following Instructions. Adherence to instructions is crucial, in tasks such as automated repair, to ensure repair quality and control costs. Figure 3(c) reveals that both GPT-4 models sometimes repeat buggy models in their initial repair iterations, despite system instructions to avoid this. GPT-4o generally shows better compliance, under all settings, with zero number of repeated buggy specifications. Interestingly, GPT-3.5-Turbo exhibited similar behavior to GPT-4o.
The evaluated LLMs exhibited varying behaviors across key factors, including instruction adherence, the number of repair iterations, and effectiveness in handling both single-line and multi-line bugs. Among all models and across all settings, GPT-4o consistently outperformed its counterparts across all measured dimensions. Notably, the dual-agent configuration generally resulted in an increased number of iterations for all models—except GPT-4o.
5.3 Results for RQ3: Characteristics of Failed Repairs